# MaxK-GNN: Extremely Fast GPU Kernel Design for Accelerating Graph Neural Networks Training

### 26 Feb 2024 v4

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; April 27-May 1, 2024; La Jolla, CA, USAbooktitle: 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’24), April 27-May 1, 2024, La Jolla, CA, USAdoi: 10.1145/3620665.3640426isbn: 979-8-4007-0385-0/24/04
@@ 1. Introduction
Graph Convolutional Networks (GCNs), a specific type of Graph Neural Networks (GNNs), have garnered significant attention in recent years due to their unparalleled capability to extract latent information from graph data (hu2020open, ; kipf2016semi, ; liu2020graphsage, ). The field of GCNs manifests in a myriad of important practical applications, including the prediction of cascading power-grid failures (liu2020guiding, ), traffic forecasting (jiang2022graph, ), recommendation systems (wu2020graph, ; ying2018graph, ), and drug discovery (bongini2021molecular, ). In the design and acceleration of GNN training, GPU platforms have become the prevalent choice. Conventional GCN acceleration processes a graph feature matrix ($X$) by multiplying it with a dense, small, weight matrix ($W$), followed by multiplying the resultant output with a highly sparse and irregular adjacency matrix ($A$) via Sparse-Dense Matrix Matrix Multiplication (SpMM) (dgl_amazon, ).
@@
We conduct extensive evaluations of MaxK-GNN system within the context of a single-GPU, full-batch, GNN training workload. We report the system training time rather than floating point operations per second (FLOPS) analysis. Experiments show that our MaxK-GNN system could approach the speedup limit in kernel design according to Amdahl’s law (gustafson1988reevaluating, ). The performance gaps between our results and the limits, e.g., 3.22$\times$/4.24$\times$ compared to 5.52$\times$/7.27$\times$ for Reddit dataset using MaxK-GNN with GraphSAGE, is from the accumulation stage of SpGEMM and dense row prefetching stage of SSpMM, which empirically are difficult to further optimize, currently.
The introduced MaxK non-linearity and kernel design are not confined to the specific framework, but exhibit compatible with other SOTA GNN training systems, including PyG (fey2019fast, ) and DGL (dgl_amazon, ). Furthermore, the adaptability of these novel constructs aligns with current methods employed in graph partitioning (chiang2019cluster, ; wan2022bns, ) and graph sampling techniques (yang2023betty, ; zeng2019graphsaint, ).
@@ 2.3. Introducing Sparsity in GNN Training
Dropout: As a well-known regularization to prevent overfitting, dropout introduces feature sparsity during training by randomly setting a fraction of input units to 0 (baldi2013understanding, ). This operation results in a form of sparsity that is highly irregular and challenging to leverage in an end-to-end hardware design.
@@
Nonlinearity for Sparsification: Nonlinear functions such as ReLU (agarap2018deep, ) introduce sparsity into the graph training. As detailed in FATReLU (kurtz2020inducing, ), adjusting the ReLU threshold can induce greater feature sparsity. Similar to other sparsity forms, this irregularity does not align with hardware characteristics, yielding limited speedup on the graph training system.
@@ 3. MaxK-GNN Dataflow
We start with introducing MaxK nonlinearity, and discuss how it can benefit GNN training system in MaxK-GNN dataflow.
@@ 3.1. MaxK Nonlinearity as a Universal Approximator
In addition, the MaxK nonlinear operator is positioned before the SpMM operator, which serves to diminish the computational and memory access overhead associated with SpMM. This nonlinearity also exhibits superior generalization ability in both transductive and inductive graph learning settings. It introduces a regularized sparsity pattern, enabling a more efficient design for hardware acceleration on GNNs.
@@
MaxK, like ReLU, is a piece-wise linear (PWL) function. We use Multilayer Perceptron (MLP) with MaxK nonlinearity to prove that it is a universal approximator. As demonstrated in Eq. 1 and as illustrated in Fig. 4(a), MaxK is applied to the feature map. The input $X$ has a size $s$, $W$ has dimensions $[s,r]$, with $r$ being the hidden dimension. MaxK maintains the maximum $k$ significant value out of $r$ while preserving the same shape. MaxK preserves the same input and output dimensions and introduces a regularized sparsity pattern, thereby facilitating the design of the supporting hardware. MaxK can approximate any continuous function $f(X)$ of $X\in\mathbb{R}^{s}$ with a sufficient number of hidden units $r$.
@@
Given any positive integers $s$ and $t$, two parameter groups $W$ and $W^{\prime}$ are determined such that $g(X)$ is expressed as a linear combination of $r$ convex PWL functions:
@@
In this equation, $g(X)$ operates as the neural network approximator and denotes the continuous PWL function with $r$ locally affine regions on $\mathbb{R}^{s}$. Thus, any continuous PWL function $g(X)$ can be represented as a linear combination of PWL functions $h(X)$ (proof can be found in (wang2004general, ; goodfellow2013maxout, )).
@@
Universal Approximator Proof: According to the Stone-Weierstrass approximation theorem (de1959stone, ), a PWL function $g(X)$ can approximate any continuous function $f(X)$ with an error $\epsilon$: $\left|f(v)-g(v)\right|<\epsilon$. The PWL function $g(X)$, as provided in Proposition 3.1, is composed of $r$ convex PWL function $h(X)$. By configuring a sufficiently large number of hidden units $r$ and appropriate $k$ values in the MaxK network $g(X)$, the desired approximation error $\epsilon$ can be achieved. Consequently, it can be concluded that a MaxK network with $r$ hidden units can provide an arbitrarily close approximation to $f(X)$ on the compact domain $C\subset\mathbb{R}^{s}$.
Refer to Fig. 4(b) for an illustration of a single-layer MaxK-based network employed for the approximation of the function $y=x^{2}$. A standard backpropagation algorithm is used to train the neural network until convergence, with the number of hidden units varied to observe the approximation error. For the MaxK nonlinearity, the top $\lceil hid/4\rceil$ elements are selected and the remainder set to 0. It can be readily observed that as the number of hidden units increases, the approximation error of the MaxK-based neural network approximator decreases. Fig. 4(c) shows the $y=x^{2}$ function approximation using ReLU nonlinearity, approximation result shows that ReLU and MaxK nonlinearity have a similar approximation performance.
@@ 4.4. Kernel Profiling
To demonstrate the effectiveness of the proposed design, We provide profiling of SpGEMM and SSpMM kernels, while the experiment setup is outlined in Sec. 5. For similicity, we use Reddit graph as an example and provide the memory system profiling result shown in Table 2. We evaluate the compute kernels by employing the Nsight Compute profiler to generate performance metrics for cuSPARSE SpMM, SpGEMM, and SSpMM kernels when executed on the Reddit graph. Table 2 reports data on the traffic between the L2 cache and global memory, as well as the hit rates for the L1 and L2 caches.
@@ 5. Evaluation
Figure 9. MaxK-GNN system evaluation for GraphSAGE (hamilton2017inductive, ), GCN (kipf2016semi, ), and GIN (xu2018gin, ) models and Reddit, ogbn-proteins, ogbn-produces, Yelp, and Flickr datasets. MaxK nonlinearity setting: $k=[2,4,8,16,32,64,96,128,192]$. We compare the end to end training speedup over DGL with cuSPARSE (dgl_amazon, ) framework and GNNAdvisor (wang2021gnnadvisor, ). spd: speedup.
@@ 5.1. Experimental Setup
Datasets. For SpGEMM & SSpMM kernel benchmark, we select popular benchmark datasets that have been extensively employed in previous studies (hamilton2017inductive, ; kipf2017semi, ; xu2019how, ; kersting2016benchmark, ; fey2019fast, ; leskovec2014snap, ; wang2021gnnadvisor, ). See Table 1 for more details.
@@ 5.3. MaxK-GNN System Training Time Evaluation
Table 5. End to end MaxK-GNN accuracy & speedup evaluation and comparison with ReLU based baseline model implemented in DGL (dgl_amazon, )
@@
The results collectively show that our MaxK-GNN could approach the speedup limit in kernel design. The performance gaps between our results and Amdahl’s law theoretical limits, i.e., 3.22$\times$/4.24$\times$ compared to 5.52$\times$/7.27$\times$ for Reddit dataset using GraphSAGE, is from the essential accumulation stage of SpGEMM and dense row prefetching stage of SSpMM, which empirically are difficult to further optimize.
@@ 6. Conclusion and Future Work
Recent advancements in MLaaS (Machine Learning as a Service) have primarily centered on accelerating inference on various platforms (xie2023accel, ; peng2023autorep, ; chen2023recontab, ; medusa, ; thorat2023advanced, ; peng2022towards, ; chen2019claims, ; peng2023lingcn, ; chen2018data, ; peng2023rrnet, ; wu2022optimizing, ; peng2022length, ; yan2023improving, ; peng2023pasnet, ; sheng2023muffin, ). In this paper, we present MaxK-GNN, a high-performance GPU training system integrating algorithm and system innovation. (i) We introduce the MaxK nonlinearity and provide a theoretical analysis of MaxK nonlinearity as a universal approximator, and present the Compressed Balanced Sparse Row (CBSR) format, designed to store the data and index of the feature matrix after nonlinearity; (ii) We design a coalescing enhanced forward computation with row-wise product-based SpGEMM Kernel using CBSR for input feature matrix fetching and strategic placement of a sparse output accumulation buffer in shared memory; (iii) We develop an optimized backward computation with outer product-based and SSpMM Kernel. Experiments show that our MaxK-GNN system could approach the limit according to Amdahl’s law. We achieve comparable accuracy to existing GNNs, but at a significantly increased speed: 3.22$\times$/4.24$\times$ speedup (vs. theoretical limits, 5.52$\times$/7.27$\times$) on Reddit.

### 19 Mar 2024 v5

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; April 27-May 1, 2024; La Jolla, CA, USAbooktitle: 29th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’24), April 27-May 1, 2024, La Jolla, CA, USAdoi: 10.1145/3620665.3640426isbn: 979-8-4007-0385-0/24/04ccs: Computer systems organization Architecturesccs: Computing methodologies Parallel computing methodologiesccs: Computing methodologies Machine learning
@@ 1. Introduction
Graph Convolutional Networks (GCNs), a specific type of Graph Neural Networks (GNNs), have garnered significant attention in recent years due to their unparalleled capability to extract latent information from graph data (hu2020open, ; kipf2016semi, ; liu2020graphsage, ). The field of GCNs manifests in a myriad of important practical applications, including the prediction of cascading power-grid failures (liu2020guiding, ), traffic forecasting (jiang2022graph, ), recommendation systems (wu2020graph, ; ying2018graph, ), and drug discovery (bongini2021molecular, ). In the design and acceleration of GNN training, GPU platforms have become the prevalent choice. Conventional GCN acceleration processes a graph feature matrix ($X$) by multiplying it with a dense small weight matrix ($W$), followed by multiplying the resultant output with a highly sparse irregular adjacency matrix ($A$) via Sparse-Dense Matrix Matrix Multiplication (SpMM) (dgl_amazon, ).
@@
We conduct extensive evaluations of MaxK-GNN system within the context of a single-GPU, full-batch, GNN training workload. We report the system training time rather than floating point operations per second (FLOPS) analysis. Experiments show that our MaxK-GNN system could approach the speedup limit according to Amdahl’s law (gustafson1988reevaluating, ). The performance gaps between our results and the limits, e.g., 3.22$\times$/4.24$\times$ compared to 5.52$\times$/7.27$\times$ for Reddit dataset using MaxK-GNN with GraphSAGE, is from the accumulation stage of SpGEMM and dense row prefetching stage of SSpMM, which empirically are difficult to further optimize.
The introduced MaxK non-linearity and kernel design are not confined to the specific framework, but exhibit compatible with other SOTA GNN training systems, including PyG (fey2019fast, ) and DGL (dgl_amazon, ). Furthermore, the adaptability of these novel constructs aligns with current methods employed in graph partitioning (chiang2019cluster, ; wan2022bns, ) and graph sampling (yang2023betty, ; zeng2019graphsaint, ).
@@ 2.3. Introducing Sparsity in GNN Training
Dropout: As a well-known regularization to prevent overfitting, dropout introduces feature sparsity during training by randomly setting a fraction of input units to 0 (baldi2013understanding, ). This operation results in a form of sparsity that is highly irregular and challenging to leverage in an system/hardware design.
@@
Nonlinearity for Sparsification: Nonlinear functions such as ReLU (agarap2018deep, ) introduce sparsity into the graph training. As detailed in FATReLU (kurtz2020inducing, ), adjusting the ReLU threshold can induce greater feature sparsity. Similar to other sparsity forms, this irregularity does not align with hardware characteristics, yielding limited speedup on the training system.
@@ 3. MaxK-GNN Dataflow
We start with introducing MaxK nonlinearity, and how it can benefit GNN training system in MaxK-GNN dataflow.
@@ 3.1. MaxK Nonlinearity as a Universal Approximator
where $W$, $b$, and $X$ represent weights, biases, and inputs, representatively. $h(X)$ denotes intermediate feature, which is a piece-wise linear (PWL) function of $X$. $r$ is the hidden dimension/number of neurons. In addition, the MaxK nonlinear operator is positioned before the SpMM operator, which serves to diminish the computational and memory access overhead associated with SpMM. This nonlinearity also exhibits generalization ability in both transductive and inductive graph learning settings. It introduces a regularized sparsity pattern, enabling a more efficient design for hardware acceleration.
@@
MaxK, like ReLU, could be represented using a PWL function. We use Multilayer Perceptron (MLP) with MaxK nonlinearity to theoretically analyze the universal approximation characteristics of MaxK nonlinearity. As demonstrated in Eq. 1 and as illustrated in Fig. 4(a), MaxK is applied to the feature map. $X$ has a size $s$. $W$ has dimensions $[s,r]$. MaxK maintains the maximum $k$ significant value out of $r$ while preserving the same shape. MaxK preserves the same input and output dimensions and introduces a regularized sparsity pattern, thereby facilitating the design of the supporting hardware. MaxK can approximate any continuous function $f(X)$ of $X\in\mathbb{R}^{s}$ with a sufficient number of hidden units $r$.
@@
Given any positive integers $s$ (input dimension) and $t$ (output dimension), two parameter groups $W$ and $W^{\prime}$ are determined such that $g(X)$ is expressed as a linear combination of $r$ convex PWL functions:
@@
$g(X)$ operates as the neural network approximator and denotes the continuous PWL function with $r$ locally affine regions on $\mathbb{R}^{s}$. Thus, $g(X)$ (could be any continuous PWL function) can be represented as a linear combination of $h(X)$ (proof can be found in (wang2004general, ; goodfellow2013maxout, )).
@@
Universal Approximator: Based on the Stone-Weierstrass approximation theorem (de1959stone, ; goodfellow2013maxout, ), a PWL function $g(X)$ can approximate any continuous function $f(X)$ with an error $\epsilon$: $\left|f(v)-g(v)\right|<\epsilon$. The PWL function $g(X)$, as provided in Proposition 3.1, is composed of $r$ number of PWL functions $h(X)$. By configuring a sufficiently large number of hidden units $r$ and appropriate $k$ values in the MaxK network $g(X)$, the desired approximation error $\epsilon$ can be achieved. Consequently, a MaxK network with $r$ hidden units could provide an arbitrarily close approximation to $f(X)$ on the compact domain $C\subset\mathbb{R}^{s}$.
Refer to Fig. 4(b) for an illustration of a single-layer MaxK-based network employed for the approximation of the function $y=x^{2}$. A standard backpropagation algorithm is used to train the neural network until convergence, with the number of hidden units varied to observe the approximation error. For the MaxK nonlinearity, the top $\lceil hid/4\rceil$ elements are selected and the remainder set to 0. It can be readily observed that as the number of hidden units increases, the approximation error of the MaxK-based neural network approximator decreases. Fig. 4(c) shows the $y=x^{2}$ function approximation using ReLU nonlinearity, results show that ReLU and MaxK nonlinearity have a similar approximation performance.
@@ 4.4. Kernel Profiling
To demonstrate the effectiveness of the proposed design, We provide profiling of SpGEMM and SSpMM kernels, while the experiment setup is outlined in Sec. 5. For similicity, we use Reddit graph as an example and provide the memory system profiling result shown in Table 2. We evaluate the compute kernels by employing the Nsight Compute profiler to generate performance metrics for cuSPARSE SpMM, SpGEMM, and SSpMM kernels when executed on the Reddit graph. Table 2 reports data on the traffic between the L2 cache and global memory, as well as the hit rates for the L1/L2 caches.
@@ 5. Evaluation
Figure 9. MaxK-GNN system evaluation for GraphSAGE (hamilton2017inductive, ), GCN (kipf2016semi, ), and GIN (xu2018gin, ) models and Reddit, ogbn-proteins, ogbn-produces, Yelp, and Flickr datasets. MaxK nonlinearity setting: $k=[2,4,8,16,32,64,96,128,192]$. We compare the MaxK-GNN training speedup over DGL with cuSPARSE (dgl_amazon, ) framework and GNNAdvisor (wang2021gnnadvisor, ). spd: speedup.
@@ 5.1. Experimental Setup
Datasets. For SpGEMM & SSpMM kernel benchmark, we select popular benchmark datasets that have been extensively employed in previous studies (hamilton2017inductive, ; kipf2017semi, ; xu2019how, ; kersting2016benchmark, ; fey2019fast, ; leskovec2014snap, ; wang2021gnnadvisor, ). See Table 1.
@@ 5.3. MaxK-GNN System Training Time Evaluation
Table 5. MaxK-GNN training accuracy & speedup evaluation and comparison with ReLU based baseline model implemented in DGL (dgl_amazon, )
@@
The results collectively show that our MaxK-GNN approaches the speedup limit. The performance gaps between our results and Amdahl’s law theoretical limits, i.e., 3.22$\times$/4.24$\times$ compared to 5.52$\times$/7.27$\times$ for Reddit dataset using GraphSAGE, is from the essential accumulation stage of SpGEMM and dense row prefetching stage of SSpMM, which empirically are difficult to further optimize.
@@ 6. Conclusion and Future Work
In this paper, we present MaxK-GNN, a high-performance GPU training system integrating algorithm and system innovation. (i) We introduce the MaxK nonlinearity and provide a theoretical analysis of MaxK nonlinearity as a universal approximator, and present the Compressed Balanced Sparse Row (CBSR) format, designed to store the data and index of the feature matrix after nonlinearity; (ii) We design a coalescing enhanced forward computation with row-wise product-based SpGEMM Kernel using CBSR for input feature matrix fetching and strategic placement of a sparse output accumulation buffer in shared memory; (iii) We develop an optimized backward computation with outer product-based and SSpMM Kernel. Experiments show that our MaxK-GNN system could approach the limit according to Amdahl’s law. We achieve comparable accuracy to existing GNNs, but at a significantly increased speed: 3.22$\times$/4.24$\times$ speedup (vs. theoretical limits, 5.52$\times$/7.27$\times$) on Reddit.